OpenStudy (anonymous):

What is the fifth term of the binomial expression?(x^2+2/8)^8 With explanation please.

OpenStudy (anonymous):

Do I just plug in 5 for x?

OpenStudy (anonymous):

Is it, 2/8?

OpenStudy (anonymous):

I was asking id that was the coefficient, but it's not.

OpenStudy (anonymous):

It's to the 8th power? Not too sure what you meant by coefficient.

jimthompson5910 (jim_thompson5910):

To generate terms of \(\Large (A+B)^{n}\), use the formula \[\Large \binom{n}{k}*(A)^{n-k}*(B)^{k}\] k is the term number that starts with k = 0 and it progresses until it hits n. In this case, we want the fifth term. So k = 4 (remember, the first term is when k = 0) Also, we have n = 8, A = x^2, B = 2/8

OpenStudy (anonymous):

I don't know how I'd impute the (n base k) in a calculator. I don't have a graphing calculator.

jimthompson5910 (jim_thompson5910):

use pascals triangle then

jimthompson5910 (jim_thompson5910):

look in the row that starts with 1, 8, ... the 5th value in that row will be equal to 8 choose 5

jimthompson5910 (jim_thompson5910):

oops sry, 8 choose 4

OpenStudy (anonymous):

70 is equal to five.

jimthompson5910 (jim_thompson5910):

so \[\Large \binom{n}{k} = \binom{8}{4} = 70\]

OpenStudy (anonymous):

So I would put 70 in for x?

jimthompson5910 (jim_thompson5910):

that's your coefficient for that formula I posted above

OpenStudy (anonymous):

Oh okay, hold on.

OpenStudy (anonymous):

Let me see if I got this.

OpenStudy (anonymous):

answer is 1792x?

jimthompson5910 (jim_thompson5910):

that doesn't sound right

jimthompson5910 (jim_thompson5910):

no it's not correct

OpenStudy (anonymous):

My options are. 1792x 1792x^4 1120x^4 1120x^3

jimthompson5910 (jim_thompson5910):

the original problem is \[\Large \left(x^2 + \frac{2}{8}\right)^8\] right?

OpenStudy (anonymous):

Correct.

jimthompson5910 (jim_thompson5910):

none of those options work, so there has to be a typo

OpenStudy (anonymous):

There's the picture @jim_thompson5910

jimthompson5910 (jim_thompson5910):

I'm getting \[\Large \frac{35}{128}x^8\] and wolfram alpha confirms it http://www.wolframalpha.com/input/?i=(x%5E2%2B2%2F8)%5E8&t=crmtb01

OpenStudy (anonymous):

Yeah, I don't understand the options either.

jimthompson5910 (jim_thompson5910):

which is why I'm thinking there's a typo